of a death-row inmate has changeed since then. She randomly selects 32 death-row inmates and finds that their mean age is 39.Please help me with this question!
Answer #1

a) The null and alternative hypothesis are:

Ho: 40.3

H 40.3

Test the hypothesis at 0.05 level of significance

b) Given

n32, 39.4, s 8.4!

The degrees of freedom are df = n-1 = 32-1 = 31

The critical value at 0.05 with 31 df is, to.05/2,31 2.04

The 95% confidence interval about the mean age is:

CI ta/2,31(-

8.4 39.4 t to.05/2,3132 = 39.4土2.04(1.484) 39.4 土3.029 - (36.371,42.429) (36.37,42.43)

The lower bound = 36.37

The upper bound = 42.43

c) Since the mean age from the earlier year is contained in the inteval,there is not sufficient evidence to conclude that the mean age has changed

Answer: option D